OpenStudy (anonymous):

Find all the solutions to the differential equation dy/dx = (tanx) (tany)? i seperated the variables and integrated to get sec^2(x)=sec^2(y) but then what?

OpenStudy (anonymous):

I think that you took derivatives.\[\int\limits_{}^{}\frac{1}{\tan y}dy=\int\limits_{}^{}\tan x dx\]I don't get secant squared anywhere.

OpenStudy (anonymous):

I changed the left side to cos y over sin y and used u substitution. I have the anti-derivative for the right side memorized, but it can actually be done in a similar way. You should get natural logs.

OpenStudy (anonymous):

ahh thnx yep i did.

OpenStudy (anonymous):

You're welcome. :)